OpenStudy (anonymous):

2m-9 7 ______ ____ 3-m 4

OpenStudy (callisto):

\[\frac{2m-9}{3-m}=\frac{7}{4}\]\[4(2m-9)=7(3-m)\]8m-36 = 21-7m Can you solve it?

OpenStudy (anonymous):

ya i think so

OpenStudy (callisto):

Basically, you just need to use cross multiplication here :)

OpenStudy (anonymous):

so 15m = 57, and m = 19/5.

OpenStudy (anonymous):

thats what i did but i guess i did something wrong in my steps and my teacher said it was wrong but i did it right with your help this time

OpenStudy (anonymous):

i got 3.8

OpenStudy (anonymous):

that's right! :D

OpenStudy (anonymous):

thanks!!!